CS209 Practicum in System Administration Quiz 2.0 April 17, 2000 1. Describe an Ethernet Switch

2. Describe an Ethernet Hub

3. Name a method for finding services provided by a computer

4. Name a common method Hackers use to acquire passwords

5. Extra Credit: How many Software Engineers does it take to screw in a lightbulb?